# Fractional binomial distributions induced by the generalized binomial theorem and their applications

### v1

@@ 2.6. The fractional multinomial distribution
We define the $(d-1)$-simplex in $\mathbb{R}^{d}$ by
@@
According to the preliminary version [FR11, Lemma 1] of the paper [FR13], we can generalize Proposition 1.1 to the multivariate cases, which is stated as follows.
Theorem 2.16 (Multivariate neo-classical inequality; see also [LCL07, Exercise 3.9]).
@@
$\begin{cases}\leq(x_{1}+x_{2}+\cdots+x_{d})^{\alpha n}&\text{if }\alpha\in(0,1],\\ \geq(x_{1}+x_{2}+\cdots+x_{d})^{\alpha n}&\text{if }\alpha\in(1,2).\\ \end{cases}$ (2.41)
@@
×$\hskip 28.45274pt\times\alpha^{d}\sum_{(j_{1},j_{2},\dots,j_{d})\in J_{d}^{(n-j_{d+1})}}\frac{\Gamma(\alpha(n-j_{d+1})+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}$
@@
Needless to say, if $\alpha=1$, the (2.41) reads as the usual multinomial theorem. As with Definition 1.3, this multivariate neo-classical inequality allows us to define the fractional analogue of the multinomial distribution.
@@
Let $\alpha>0$, $n\in\mathbb{N}$ and $\bm{x}=(x_{1},x_{2},\dots,x_{d})\in\Delta_{d-1}$. Then, a probability measure $\mu_{\alpha,\bm{x}}^{(n)}$ on $J_{d}^{(n)}$ is called the $\alpha$-fractional multinomial distribution if
$\mu_{\alpha,\bm{x}}^{(n)}(\{\bm{j}\})=\frac{\alpha^{d}}{Z_{\alpha,\bm{x}}^{(n)}}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}$ (2.42)
for $\bm{j}\in J_{d}^{(n)}$, where $Z_{\alpha,\bm{x}}^{(n)}$ is the normalized constant given by
@@
for $d>2$. Obtaining such a formula explicitly might require a multivariate version of the fractional Taylor series, which although seeming rather difficult might nevertheless be worth investigating further.
3. Limit theorems for the fractional binomial distribution
3.1. The weak law of large numbers
First, we prove the weak law of large numbers for the $\alpha$-fractional binomial distribution $\mu_{\alpha,x}^{(n)}$.
@@ 3.1. The weak law of large numbers
$\leq\frac{1}{\varepsilon^{2}}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}\right|=O\left(\frac{1}{n}\right)$
@@
3.2. The central limit theorem
We show the central limit theorem for the $\alpha$-fractional binomial distribution $\mu_{\alpha,x}^{(n)}$, which is stated as follows.
@@ 3.2. The central limit theorem
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. Let $\widetilde{\mu}_{\alpha,x}^{(n)}$ denote the probability law of the random variable
@@
$\hskip 28.45274pt+\frac{1}{Z_{\alpha,x}^{(n)}}\sum_{\omega\in K_{\alpha}\setminus\{1\}}\left(1-x+x\omega\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\right)\right)^{\alpha n}$
$\hskip 28.45274pt-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha\sin\alpha\pi}{\pi}\exp\left(\frac{i\xi}{\sqrt{v_{n}}}\right)\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
×$\hskip 28.45274pt\times\Bigg{\{}\frac{x^{\alpha}(1-x)^{\alpha(n+1)}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi/\sqrt{v_{n}}-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi/\sqrt{v_{n}}+\alpha\pi)}\}}$
$\hskip 28.45274pt+\frac{x^{\alpha(n+1)}(1-x)^{\alpha}e^{i\xi(n-2)/\sqrt{v_{n}}}}{\{(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}+\alpha\pi)}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}-\alpha\pi)}-(tx)^{\alpha}\}}\Bigg{\}}\,\mathrm{d}t\Bigg{]}.$ (3.2)
@@
=$=\left\{(1-x)\exp\left(-\frac{i\xi m_{n}}{\alpha n\sqrt{v_{n}}}\right)+x\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\left(1-\frac{m_{n}}{n}\right)\right)\right\}^{\alpha n}$
@@
$\hskip 28.45274pt+x\left\{1+\frac{i\xi}{\alpha\sqrt{v_{n}}}\left(1-\frac{m_{n}}{n}\right)-\frac{\xi^{2}}{2(\alpha\sqrt{v_{n}})^{2}}\left(1-\frac{m_{n}}{n}\right)^{2}+O(n^{-3/2})\right\}\Bigg{]}^{\alpha n}$
@@
$\hskip 28.45274pt+x\left\{1+\frac{i\xi(1-x)}{\sqrt{\alpha nx(1-x)}}-\frac{\xi^{2}(1-x)^{2}}{2\alpha nx(1-x)}+O(n^{-3/2})\right\}\Bigg{]}^{\alpha n}$
=$=\left\{1-\frac{\xi^{2}}{2\alpha n}+O(n^{-3/2})\right\}^{\alpha n}\to\exp\left(-\frac{\xi^{2}}{2}\right)$ (3.3)
@@
as $n\to\infty$, (2.13) implies that the third term of the right-hand side of (3.2) converges to zero as $n\to\infty$. Therefore, (3.2), (3.3), and (3.4) lead to
@@
Because the right-hand side of the above is nothing but the characteristic function of the normal distribution $N(0,1)$, Lévy’s continuity theorem (cf. [Kle08, Theorem 15.23]) concludes the desired weak convergence. ∎
3.3. The law of small numbers
@@
Let $\alpha>0$ and $\lambda>0$. The $\alpha$-fractional Poisson distribution with parameter $\lambda$ is a probability measure $\nu_{\alpha,\lambda}$ on $\mathbb{N}_{0}$ defined as
$\nu_{\alpha,\lambda}(\{j\})=\frac{1}{E_{\alpha}(\lambda^{\alpha})}\frac{\lambda^{\alpha j}}{\Gamma(\alpha j+1)},\qquad j\in\mathbb{N}_{0}.$
The $1$-fractional Poisson distribution coincides with the Poisson distribution. Some basic properties of this fractional Poisson distribution will be summarized in the subsequent section. We note that $\nu_{\alpha,\lambda}$ is a special case of generalized fractional Poisson distributions defined via the generalized Mittag-Leffler function $E_{\alpha,\beta}(z)$ given by (3.9) below. See [Her16, CO17] for some properties as well as its definition.
Here, we show that $\nu_{\alpha,\lambda}$ is captured through the law of small numbers for the fractional binomial distribution.
@@
Let $\alpha>0$, $\lambda>0$, and $x\in(0,1)$. Suppose that two sequences $\{\alpha_{n}\}_{n=1}^{\infty}\subset(0,\infty)$ and $\{x_{n}\}_{n=1}^{\infty}\subset(0,1)$ satisfy that $\alpha_{n}\to\alpha$ and $\alpha_{n}nx_{n}\to\lambda$ as $n\to\infty$. Then, $\mu_{\alpha_{n},x_{n}}^{(n)}$ converges weakly to $\nu_{\alpha,\lambda}$ as $n\to\infty$.
@@
Because $\alpha_{n}\to\alpha>0$ and $\alpha_{n}nx_{n}\to\lambda>0$, there exist four constants $C_{3,1},C_{3,2},C_{3,3},C_{3,4}>0$ such that
@@
(0(0( 0 $\leq)\,b_{\alpha_{n},x_{n}}^{(n)}(j)$
$=\frac{\alpha_{n}\Gamma(\alpha_{n}n+1)}{\Gamma(\alpha_{n}(n-j)+1)\Gamma(\alpha_{n}j+1)}x_{n}^{\alpha_{n}j}(1-x_{n})^{\alpha_{n}(n-j)}$
@@
$(0\leq)\,b_{\alpha_{n},x_{n}}^{(n)}(j)$ =$=\begin{cases}\displaystyle C_{3,2}e^{-C_{3,1}C_{3,3}}&\text{if }j=0,\vspace{2mm}\\ \displaystyle\frac{C_{3,2}C_{3,6}(C_{3,2}C_{3,4}\vee 1)^{C_{3,2}j}}{C_{3,5}\Gamma(C_{3,1}j+1)}(j+1)^{1/2}&\text{if }j=1,2,\dots,n-1,\vspace{2mm}\\ \displaystyle C_{3,2}\left(\frac{C_{3,4}}{j}\right)^{C_{3,1}j}&\text{if }j=n,\\ 0&\text{if }j>n\end{cases}$
@@
3.4. Some properties of the fractional Poisson distribution
@@
Let $S_{\alpha,\lambda}$ be a random variable whose probability law is $\nu_{\alpha,\lambda}$. Then, we have
@@
$\hskip 28.45274pt+\lambda^{\alpha}E_{\alpha}(\lambda^{\alpha})\big{\{}E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\big{\}}-\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})^{2}\Big{]},$
@@
$\frac{\mathrm{d}^{2}}{\mathrm{d}\xi^{2}}\varphi_{\alpha,\lambda}(\xi)$ $=i\frac{\lambda^{\alpha}}{\alpha E_{\alpha}(\lambda^{\alpha})}\Bigg{\{}ie^{i\xi}E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})$
$\hskip 28.45274pt+i\lambda^{\alpha}e^{2i\xi}\left(\frac{1}{\alpha}E_{\alpha,2\alpha-1}(\lambda^{\alpha}e^{i\xi})-\frac{\alpha-1}{\alpha}E_{\alpha,2\alpha}(\lambda^{\alpha}e^{i\xi})\right)\Bigg{\}}$
$=-\frac{\lambda^{\alpha}e^{i\xi}}{\alpha E_{\alpha}(\lambda^{\alpha})}\Bigg{\{}E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})$
$\hskip 28.45274pt+\lambda^{\alpha}e^{i\xi}\left(\frac{1}{\alpha}E_{\alpha,2\alpha-1}(\lambda^{\alpha}e^{i\xi})-\frac{\alpha-1}{\alpha}E_{\alpha,2\alpha}(\lambda^{\alpha}e^{i\xi})\right)\Bigg{\}}.$ (3.13)
@@
$\hskip 28.45274pt+\sum_{j=0}^{\infty}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)^{m}\frac{\alpha}{\lambda}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)\nu_{\alpha,\lambda}(\{j\})$
@@
$\frac{\mathrm{d}}{\mathrm{d}\lambda}\left(\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)$
$=\frac{\lambda^{\alpha-1}}{\alpha E_{\alpha}(\lambda^{\alpha})^{2}}\Big{[}\alpha E_{\alpha}(\lambda^{\alpha})E_{\alpha,\alpha}(\lambda^{\alpha})$
$\hskip 28.45274pt+\lambda^{\alpha}E_{\alpha}(\lambda^{\alpha})\big{\{}E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\big{\}}-\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})^{2}\Big{]}$
@@
4. The fractional Bernstein operator and limit theorems for its iterates
4.1. The fractional Bernstein operator and its approximating property
Let $\alpha>0$. In this section, we show that the fractional Bernstein operator $B_{\alpha,n}$ introduced in Definition 1.8 uniformly approximates all continuous functions on $[0,1]$, similarly to Proposition 1.5.

### v3

@@ 2.6 The fractional multinomial distribution
We define the (d1)(d-1)-simplex in $\mathbb{R}^{d}$ by
@@
According to the preliminary version [8, Lemma 1] of the paper [9], we can generalize the neo-classical inequality presented in [12, Theorem 1.2] to the multivariate cases, which is stated as follows.
Theorem 2.16 (Multivariate neo-classical inequality; see also [23, Exercise 3.9]).
@@
[…]
@@
$\quad\times\alpha^{d}\sum_{(j_{1},j_{2},\dots,j_{d})\in J_{d}^{(n-j_{d+1})}}\frac{\Gamma(\alpha(n-j_{d+1})+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}$
@@
Needless to say, if $\alpha=1$, the (2.41) reads as usual multinomial theorem. As with Definition 1.2, this multivariate neo-classical inequality allows us to define the fractional analogue of the multinomial distribution.
@@
Let $\alpha>0$, $n\in\mathbb{N}$ and $\bm{x}=(x_{1},x_{2},\dots,x_{d})\in\Delta_{d-1}$. Then, a probability measure $\mu_{\alpha,\bm{x}}^{(n)}$ is called the $\alpha$-fractional multinomial distribution if
$\mu_{\alpha,\bm{x}}^{(n)}(\mathrm{d}\bm{z}):=\frac{\alpha^{d}}{Z_{\alpha,\bm{x}}^{(n)}}\sum_{\bm{j}\in J_{d}^{(n)}}\frac{\Gamma(\alpha n+1)}{\Gamma(\alpha j_{1}+1)\Gamma(\alpha j_{2}+1)\cdots\Gamma(\alpha j_{d}+1)}x_{1}^{\alpha j_{1}}x_{2}^{\alpha j_{2}}\cdots x_{d}^{\alpha j_{d}}\delta_{\bm{j}}(\mathrm{d}\bm{z}),$ (2.42)
where $Z_{\alpha,\bm{x}}^{(n)}$ is the normalized constant given by
@@
for $d>2$. Obtaining such a formula explicitly might require a multivariate version of the fractional Taylor series, which, although it appears rather difficult, may nevertheless be worth further investigation.
3 Limit theorems for the fractional binomial distribution
3.1 The weak law of large numbers
First, we prove the weak law of large numbers for the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$.
@@ 3.1 The weak law of large numbers
$\leq\frac{1}{\varepsilon^{2}}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}\right|$
$=O\left(\frac{1}{n}\right)$
@@
3.2 The central limit theorem
We prove the central limit theorem for the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$, which is stated as follows.
@@ 3.2 The central limit theorem
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. Let $\widetilde{\mu}_{\alpha,x}^{(n)}$ denote the law of the random variable
@@
$\quad+\frac{1}{Z_{\alpha,x}^{(n)}}\sum_{\omega\in K_{\alpha}\setminus\{1\}}\left(1-x+x\omega\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\right)\right)^{\alpha n}$
$\quad-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha\sin\alpha\pi}{\pi}\exp\left(\frac{i\xi}{\sqrt{v_{n}}}\right)\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
$\quad\times\Bigg\{\frac{x^{\alpha}(1-x)^{\alpha(n+1)}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi/\sqrt{v_{n}}-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi/\sqrt{v_{n}}+\alpha\pi)}\}}$
$\qquad+\frac{x^{\alpha(n+1)}(1-x)^{\alpha}e^{i\xi(n-2)/\sqrt{v_{n}}}}{\{(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}+\alpha\pi)}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{-i(\xi/\sqrt{v_{n}}-\alpha\pi)}-(tx)^{\alpha}\}}\Bigg\}\,\mathrm{d}t\Bigg].$ (3.2)
@@
$=\left\{(1-x)\exp\left(-\frac{i\xi m_{n}}{\alpha n\sqrt{v_{n}}}\right)+x\exp\left(\frac{i\xi}{\alpha\sqrt{v_{n}}}\left(1-\frac{m_{n}}{n}\right)\right)\right\}^{\alpha n}$
@@
$\quad+x\left\{1+\frac{i\xi}{\alpha\sqrt{v_{n}}}\left(1-\frac{m_{n}}{n}\right)-\frac{\xi^{2}}{2(\alpha\sqrt{v_{n}})^{2}}\left(1-\frac{m_{n}}{n}\right)^{2}+O(n^{-3/2})\right\}\Bigg]^{\alpha n}$
@@
$\quad+x\left\{1+\frac{i\xi(1-x)}{\sqrt{\alpha nx(1-x)}}-\frac{\xi^{2}(1-x)^{2}}{2\alpha nx(1-x)}+O(n^{-3/2})\right\}\Bigg]^{\alpha n}$
$=\left\{1-\frac{\xi^{2}}{2\alpha n}+O(n^{-3/2})\right\}^{\alpha n}\to\exp\left(-\frac{\xi^{2}}{2}\right)$ (3.3)
@@
as $n\to\infty$, Proposition 2.6 (1) implies that the third term of the right-hand side of (3.2) converges to zero as $n\to\infty$. Therefore, (3.2), (3.3), and (3.4) lead to
@@
Because the right-hand side of the above is nothing but the characteristic function of the normal distribution $N(0,1)$, Lévy’s continuity theorem (cf. [17, Theorem 15.23]) concludes the desired weak convergence. ∎
By combining Corollary 2.8 and Theorem 3.2, we also have the following.
Corollary 3.3.
In the notation of Theorem 3.2, the law of $\frac{S_{\alpha,x}^{(n)}-nx}{\sqrt{nx(1-x)/\alpha}}$ converges weakly to $N(0,1)$ as $n\to\infty$.
3.3 The law of small numbers
@@
Let $\alpha>0$ and $\lambda>0$. The $\alpha$-fractional Poisson distribution with parameter $\lambda$ is a probability measure $\nu_{\alpha,\lambda}$ defined as
$\nu_{\alpha,\lambda}(\mathrm{d}z):=\frac{1}{E_{\alpha}(\lambda^{\alpha})}\sum_{j=0}^{\infty}\frac{\lambda^{\alpha j}}{\Gamma(\alpha j+1)}\delta_{j}(\mathrm{d}z).$
The $1$-fractional Poisson distribution coincides with the Poisson distribution. Some basic properties of this fractional Poisson distribution will be summarized in the subsequent section. We note that $\nu_{\alpha,\lambda}$ is a special case of generalized fractional Poisson distributions defined via the generalized Mittag-Leffler function $E_{\alpha,\beta}(z)$ given by (3.9) below. See [13, 6] for some properties as well as its definition. We also refer to [19] for the study of non-Markovian fractional Poisson processes based on Riemann–Liouville type fractional calculus, and [21] for some generalizations of Bernoulli processes having long-range dependence and their convergences to fractional Poisson processes, both of which are different concepts from ours. We also mention that the Mittag-Leffler function has already been used to introduce some probability distributions in different situations. See [26, 27] for the notion of the Mittag-Leffler distribution, which is, however, far from the Poisson distribution.
Here, we prove that $\nu_{\alpha,\lambda}$ is captured through the law of small numbers for the fractional binomial distribution.
@@
Let $\alpha>0$ and $\lambda>0$. Suppose that two sequences $\{\alpha_{n}\}_{n=1}^{\infty}\subset(0,\infty)$ and $\{x_{n}\}_{n=1}^{\infty}\subset(0,1)$ satisfy that $\alpha_{n}\to\alpha$ and $nx_{n}\to\lambda/\alpha$ as $n\to\infty$. Then, $\mu_{\alpha_{n},x_{n}}^{(n)}$ converges weakly to $\nu_{\alpha,\lambda}$ as $n\to\infty$.
@@
Because $\alpha_{n}\to\alpha>0$ and $nx_{n}\to\lambda/\alpha>0$, there exist four constants $C_{3,1},C_{3,2},C_{3,3},C_{3,4}>0$ such that
@@
$(0\leq)\,b_{\alpha_{n},x_{n}}^{(n)}(j)$ $=\frac{\alpha_{n}\Gamma(\alpha_{n}n+1)}{\Gamma(\alpha_{n}(n-j)+1)\Gamma(\alpha_{n}j+1)}x_{n}^{\alpha_{n}j}(1-x_{n})^{\alpha_{n}(n-j)}$
@@
$(0\leq)\,b_{\alpha_{n},x_{n}}^{(n)}(j)$ $=\begin{cases}\displaystyle C_{3,2}e^{-C_{3,1}C_{3,3}}&\text{if }j=0,\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\frac{C_{3,2}C_{3,6}(C_{3,2}C_{3,4}\vee 1)^{C_{3,2}j}}{C_{3,5}\Gamma(C_{3,1}j+1)}(j+1)^{1/2}&\text{if }j=1,2,\dots,n-1,\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle C_{3,2}\left(\frac{C_{3,4}}{j}\right)^{C_{3,1}j}&\text{if }j=n,\\ 0&\text{if }j>n\end{cases}$
@@
3.4 Some properties of the fractional Poisson distribution
@@
Let $S_{\alpha,\lambda}$ be a random variable whose law is $\nu_{\alpha,\lambda}$. Then, we have
@@
$\quad+\lambda^{\alpha}E_{\alpha}(\lambda^{\alpha})\big\{E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\big\}-\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})^{2}\Big],$
@@
$\frac{\mathrm{d}^{2}}{\mathrm{d}\xi^{2}}\varphi_{\alpha,\lambda}(\xi)$
$=i\frac{\lambda^{\alpha}}{\alpha E_{\alpha}(\lambda^{\alpha})}\Bigg\{ie^{i\xi}E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})+i\lambda^{\alpha}e^{2i\xi}\left(\frac{1}{\alpha}E_{\alpha,2\alpha-1}(\lambda^{\alpha}e^{i\xi})-\frac{\alpha-1}{\alpha}E_{\alpha,2\alpha}(\lambda^{\alpha}e^{i\xi})\right)\Bigg\}$
$=-\frac{\lambda^{\alpha}e^{i\xi}}{\alpha E_{\alpha}(\lambda^{\alpha})}\Bigg\{E_{\alpha,\alpha}(\lambda^{\alpha}e^{i\xi})+\lambda^{\alpha}e^{i\xi}\left(\frac{1}{\alpha}E_{\alpha,2\alpha-1}(\lambda^{\alpha}e^{i\xi})-\frac{\alpha-1}{\alpha}E_{\alpha,2\alpha}(\lambda^{\alpha}e^{i\xi})\right)\Bigg\}.$ (3.13)
@@
$\quad+\sum_{j=0}^{\infty}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)^{m}\frac{\alpha}{\lambda}\left(j-\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)\nu_{\alpha,\lambda}(\{j\})$
@@
$\frac{\mathrm{d}}{\mathrm{d}\lambda}\left(\frac{\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})}{\alpha E_{\alpha}(\lambda^{\alpha})}\right)$ $=\frac{\lambda^{\alpha-1}}{\alpha E_{\alpha}(\lambda^{\alpha})^{2}}\Big[\alpha E_{\alpha}(\lambda^{\alpha})E_{\alpha,\alpha}(\lambda^{\alpha})$
$\quad+\lambda^{\alpha}E_{\alpha}(\lambda^{\alpha})\big\{E_{\alpha,2\alpha-1}(\lambda^{\alpha})-(\alpha-1)E_{\alpha,2\alpha}(\lambda^{\alpha})\big\}-\lambda^{\alpha}E_{\alpha,\alpha}(\lambda^{\alpha})^{2}\Big]$
@@
4 The fractional Bernstein operator and limit theorems for its iterates
4.1 The fractional Bernstein operator and its approximating property
Let $\alpha>0$. In this section, we prove that the fractional Bernstein operator $B_{\alpha,n}$ introduced in Definition 1.3 uniformly approximates all continuous functions on $[0,1]$, similarly to the classical case.